Clara earned $360 in simple interest on a 12-month bank certificate of deposit. If the interest rate had been 3% higher, she would have earned the same amount of money by depositing $1,000 less. How much money did Clara deposit, and what was her interest rate?
Answer by stanbon(75887)

Clara earned $360 in simple interest on a 12-month bank certificate of deposit. If the interest rate had been 3% higher, she would have earned the same amount of money by depositing $1,000 less. How much money did Clara deposit, and what was her interest rate?
----------------------
Interest=pricipal*rate*time
EQUATIONS:
1st: 360=Pr
2nd: 360=(P-1000)(1.03r)
----------
Solve 1st for r:
r=360/P
Substitute into 2nd to solve for P:
360=(P-1000)(1.03)(360/P)
360=(P-1000)(370/P)
360=370-370000/P
-10=-370000/P
P=$37,000 (amount of money she deposited)
-------------
Substitute to find r:
r=360/37000
r=0.9% (her interest rate)
--------------------
Cheers,
Stan H.